# Trilinear characterizations of the Fourier extension conjecture on the paraboloid in three dimensions

### v3

@@ Abstract.
for all $\nu$-disjoint triples $\left(U_{1},U_{2},U_{3}\right)$, all $\varepsilon>0$, $s\in\mathbb{N}$ and $s_{1},s_{2},s_{3}$ arbitrarily close to $s$.
@@ 3.1. Initial setup and statement of the main Alpert characterization
for all $r\in\mathbb{N}$, all $\nu$-disjoint triples $\left(U_{1},U_{2},U_{3}\right)$, all smooth Alpert pseudoprojections $\mathsf{Q}_{s_{k},U_{k}}^{\eta}$ with moment vanishing parameter $\kappa$ and with $\frac{r}{1+\delta}<s_{1}\leq s_{2}\leq s_{3}<\frac{r}{1-\delta}$, and all $f_{1},f_{2},f_{3}\in L^{\infty}$.
@@
The disjoint smooth Alpert trilinear inequality $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa,\delta}\left(\otimes L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ represents the weakest formulation of the Fourier extension conjecture that the authors could find to date, requiring only smooth Alpert projections of bounded functions at scales near $s$ in a disjoint trilinear inequality, integration over the corresponding ‘small’ annulus $A\left(0,2^{s}\right)$ of resonance, and permitting the familiar small $\varepsilon$-power growth in $s$.
@@ 3.4. Case 2: $s_{2}$ is small
3.5. Case 3: $s_{1}$ is small, but not $s_{2}$
Proof continued.
Here we show that (3.6) holds in the case when $s_{1}$ and $s_{2}$ are slightly separated, i.e.
$s_{1}<\frac{r}{1+2\delta}<\frac{r}{1+\delta}<s_{2}\leq s_{3}<\frac{r}{1-\delta}.$
For this we use the smoothness of $\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}$ together with the moment vanishing of $\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}$ and the $\nu$-separation of $U_{1}$ and $U_{2}$ as follows. Start with
(3.13) $\left|\mathcal{E}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\ \mathcal{E}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|=\left|\left[\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\right)\ast\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\right)\right]^{\wedge}\left(\xi\right)\right|$
$\lesssim\left\|\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\right)\ast\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\right)\right\|_{L^{1}\left(\mathbb{R}^{3}\right)}$
$\leq\left\|\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\right)\ast\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\right)\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\left|\Phi\left(U_{1}\right)+\Phi\left(U_{2}\right)\right|,$
where the $L^{1}\left(\mathbb{R}^{3}\right)$ norm is applied to the absolutely continuous measure $\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\right)\ast\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\right)$, and the $L^{\infty}\left(\mathbb{R}^{3}\right)$ norm is applied to the density $a\rightarrow\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\right)\ast\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\right)\left(a\right)$ as given in (3.9), which is smoothly adapted to scale $\frac{1}{\nu}2^{-s_{1}}$ by Lemma 11. We now claim that from (3.8), (3.9) and (3.10) we have
(3.14) $\left\|\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\right)\ast\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\right)\right\|_{L^{\infty}}\lesssim\frac{1}{\nu^{\kappa}}2^{-\kappa\left(s_{2}-s_{1}\right)}2^{\frac{2}{p}s_{1}}2^{\frac{2}{p}s_{2}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}.$
Indeed, with $\mu^{1}\equiv\Phi_{\ast}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}$ and $\mu^{2}\equiv\Phi_{\ast}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}$, we have from (3.7) that
$\mu^{1}\ast\mu^{2}\left(a\right)$ $=\iint_{u\in U_{2}}\left\{\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\left(u+a^{\prime}\right)\beta_{a}\left(u,u+a^{\prime}\right)\right\}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\left(u\right)du$
$=\iint_{u\in U_{2}}H_{a}\left(u\right)\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\left(u\right)du,$
where $H_{a}\left(u\right)$ is a smooth function of $u$ which is smoothly adapted to scale $\frac{1}{\nu}2^{-s_{1}}$. Since $\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}$ has moment vanishing at scale $2^{-s_{2}}$, we obtain (3.14) by Taylor’s formula as in Case 1.
From (3.10) we obtain that for $1<p<\infty$,
$\left\|\mathcal{E}\mathsf{Q}_{\kappa,s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ $\lesssim\left\|\mathsf{Q}_{\kappa,s_{3},U_{3}}^{\eta}f_{3}\right\|_{L^{1}\left(U_{3}\right)}=\left\|\sum_{I\in\mathcal{G}_{s_{3}}\left[U_{3}\right]}\ell\left(I\right)^{-\frac{2}{p}}\left\|f_{3}\right\|_{L^{p}}\mathbf{1}_{\left(1+\eta\ell\left(I\right)\right)I}\left(x\right)\right\|_{L^{1}\left(U_{3}\right)}$
$\leq 2^{\frac{2}{p}s_{3}}\sum_{I\in\mathcal{G}_{s_{3}}\left[U_{3}\right]}\left\|\mathbf{1}_{\left(1+\eta\ell\left(I\right)\right)I}\left(x\right)\right\|_{L^{1}\left(U_{3}\right)}\left\|f_{3}\right\|_{L^{p}}\lesssim 2^{\frac{2}{p}s_{3}}\left\|f_{3}\right\|_{L^{p}}\ .$
Now we can prove Case 3 using
$s_{2}-s_{1}\geq\frac{r}{1+\delta}-\frac{r}{1+2\delta}=\frac{\delta}{\left(1+\delta\right)\left(1+2\delta\right)}r\geq\frac{\delta}{6}r.$
Indeed, (3.13) and (3.14) give
$\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{\kappa,s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\lesssim\left(\int_{A\left(0,2^{r}\right)}\left(\left\|\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\right)\ast\left(\Phi_{\ast}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\right)\right\|_{L^{p}\left(\mathbb{R}^{3}\right)}\left|\Phi\left(U_{1}\right)+\Phi\left(U_{2}\right)\right|\ 2^{\frac{2}{p}s_{3}}\left\|f_{3}\right\|_{L^{p}\left(U_{3}\right)}\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\lesssim\left|A\left(0,2^{r}\right)\right|^{\frac{3}{q}}\frac{1}{\nu^{\kappa}}2^{-\kappa\left(s_{2}-s_{1}\right)}2^{\frac{2}{p}\left(\frac{3r}{1-\delta}\right)}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}\left(U_{3}\right)}\ \ \ \ \ \text{(since }s_{1}+s_{2}+s_{3}<\frac{3r}{1-\delta}\text{)}$
$\leq\frac{1}{\nu^{\kappa}}2^{\left(\frac{9}{q}+\frac{6}{p\left(1-\delta\right)}\right)r}2^{-\kappa\frac{\delta}{6}r}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}\left(U_{3}\right)},$
where $2^{\left(\frac{9}{q}+\frac{6}{p\left(1-\delta\right)}\right)r}2^{-\kappa\frac{\delta}{6}r}\leq 1$ for $\kappa>\left(\frac{54}{q}+\frac{36}{p\left(1-\delta\right)}\right)\frac{1}{\delta}$. Thus if we take $\kappa>\frac{20}{\delta}$, $0<\delta<\frac{1}{2}$, $q>3$ and $p>36$, then we have shown that
(3.15) $\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{\kappa,s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{\kappa,s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{\kappa,s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
$\leq\frac{1}{\nu^{\kappa}}\left\|f_{1}\right\|_{L^{p}}\left\|f_{2}\right\|_{L^{p}}\left\|f_{3}\right\|_{L^{p}\left(U_{3}\right)},$
holds in Case 3, i.e.
$s_{1}<\frac{r}{1+2\delta}<\frac{r}{1+\delta}<s_{2}\leq s_{3}<\frac{r}{1-\delta},$
provided $p>36$. ∎
@@ 3.6. Wrapup of the proof
where the restriction to $I_{k}\in\mathcal{G}\left[U_{k}\right]$ can be made by adapting the reduction argument in the proof of [Saw7, Lemma 2 on page 5]. Next we obtain from (3.11), (3.12) and (3.15), that the last line above is at most,
$\left\{\sum_{\frac{r}{1+2\delta}\leq s_{1}\leq s_{2},s_{3}\leq\frac{r}{1-\delta}}+\sum_{s_{1}\leq\frac{r}{1+2\delta},\frac{r}{1+\delta}\leq s_{2},s_{3}\leq\frac{r}{1-\delta}}+\sum_{s_{1}\leq s_{2}\leq s_{3}\text{ and }\frac{r}{1-\delta}\leq s_{3}}+\sum_{s_{1}\leq s_{2}\leq s_{3}\leq\frac{r}{1-\delta}\text{ and }s_{1}\leq s_{2}\leq\frac{r}{1+\delta}}\right\}$
@@
$\lesssim\sum_{\frac{r}{1+2\delta}\leq s_{1}\leq s_{2},s_{3}\leq\frac{r}{1-\delta}}\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
@@
which is $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$, where we have used $R^{\varepsilon}+C_{\nu}\leq C_{\nu}R^{\varepsilon}$ for $R^{\varepsilon},C_{\nu}\geq 1$. By Theorem 3, this is equivalent to the Fourier extension conjecture, and this completes the proof of Theorem 10, since we may trivially replace the exponent range $\frac{r}{1+2\delta}\leq s_{1}\leq s_{2},s_{3}\leq\frac{r}{1-\delta}$ with $\frac{r}{1+\delta}\leq s_{1}\leq s_{2},s_{3}\leq\frac{r}{1-\delta}$, as $\delta>0$ is arbitrary. ∎
@@ 4. Appendix: Application to the probabilistic Fourier extension theorem
taken over all $s_{k}\in\left(\left(1-\delta\right)s,\left(1+\delta\right)s\right)$, all $f_{k}\in L^{q}\left(U_{k}\right)$, and all triples $\left(U_{1},U_{2},U_{3}\right)\subset U^{3}$ that satisfy the weak $\nu$-disjoint condition,
@@
It should be possible to prove this square function variant by tracing through the proof of Theorem 10 above and making modifications for the square function. See for example [RiSa2], where these modifications are carried out for ‘father’ wavelets $\varphi_{I}$ smoothly adapted to a square $I$, and their projections $\bigtriangleup_{I}f\equiv\left\langle f,\varphi_{I}\right\rangle\varphi_{I}$. However, an additional difficulty arises for smooth Alpert wavelets $h_{I;\kappa}^{\eta}$, since the pseudoprojections $\bigtriangleup_{I;\kappa}^{\eta}f\equiv\left\langle S_{\kappa,\eta}^{-1}f,h_{I;\kappa}\varphi_{I}\right\rangle h_{I;\kappa}^{\eta}$ involve an operator $S_{\kappa,\eta}^{-1}$ that is bounded only on $L^{p}$ for $1<p<\infty$, and not on $L^{\infty}$. In this case, parabolic rescalings must be performed with $f\in L^{p}\left(U\right)$, which introduces an additional growth factor $C_{p}2^{\frac{2}{p}s}$, that turns out to be harmless because one can take $p$ arbitrarily large. As mentioned earlier, the proof of this conjecture will be addressed in a future paper.

### v4

@@ Abstract.
for all $\nu$-disjoint triples $\left(U_{1},U_{2},U_{3}\right)$, all $\varepsilon>0$, $s\in\mathbb{N}$ and $s_{1}\leq s_{2}\leq s_{3}$ with $s_{2},s_{3}$ arbitrarily close to $s$.
@@ 3.1. Initial setup and statement of the main Alpert characterization
for all $r\in\mathbb{N}$, all $\nu$-disjoint triples $\left(U_{1},U_{2},U_{3}\right)$, all smooth Alpert pseudoprojections $\mathsf{Q}_{s_{k},U_{k}}^{\eta}$ with moment vanishing parameter $\kappa$ and with $s_{1}\leq s_{2}$ and $\frac{r}{1+\delta}<s_{2}\leq s_{3}<\frac{r}{1-\delta}$, and all $f_{1},f_{2},f_{3}\in L^{\infty}$.
@@
The disjoint smooth Alpert trilinear inequality $\mathcal{A}_{\mathop{\rm disj}\nu}^{\kappa,\delta}\left(\otimes L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$ represents the weakest formulation of the Fourier extension conjecture that the authors could find to date, requiring only smooth Alpert projections of bounded functions with the two largest scales near $s$ in a disjoint trilinear inequality, integration over the corresponding ‘small’ annulus $A\left(0,2^{s}\right)$ of resonance, and permitting the familiar small $\varepsilon$-power growth in $s$.
@@ 3.5. Wrapup of the proof
where the restriction to $I_{k}\in\mathcal{G}\left[U_{k}\right]$ can be made by adapting the reduction argument in the proof of [Saw7, Lemma 2 on page 5]. Next we obtain from (3.11) and (3.12), that the last line above is at most,
$\left\{\sum_{s_{1}\leq s_{2}\leq s_{3}\text{ and }\frac{r}{1+\delta}\leq s_{2},s_{3}\leq\frac{r}{1-\delta}}+\sum_{s_{1}\leq s_{2}\leq s_{3}\text{ and }\frac{r}{1-\delta}\leq s_{3}}+\sum_{s_{1}\leq s_{2}\leq s_{3}\leq\frac{r}{1-\delta}\text{ and }s_{1}\leq s_{2}\leq\frac{r}{1+\delta}}\right\}$
@@
$\lesssim\sum_{s_{1}\leq s_{2}\leq s_{3}\text{ and }\frac{r}{1+\delta}\leq s_{2},s_{3}\leq\frac{r}{1-\delta}}\left(\int_{A\left(0,2^{r}\right)}\left(\left|\mathcal{E}\mathsf{Q}_{s_{1},U_{1}}^{\eta}f_{1}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{2},U_{2}}^{\eta}f_{2}\left(\xi\right)\right|\ \left|\mathcal{E}\mathsf{Q}_{s_{3},U_{3}}^{\eta}f_{3}\left(\xi\right)\right|\right)^{\frac{q}{3}}\ d\xi\right)^{\frac{3}{q}}$
@@
which is $\mathcal{E}_{\mathop{\rm disj}\nu}\left(\otimes_{3}L^{\infty}\rightarrow L^{\frac{q}{3}};\varepsilon\right)$, where we have used $R^{\varepsilon}+C_{\nu}\leq C_{\nu}R^{\varepsilon}$ for $R^{\varepsilon},C_{\nu}\geq 1$. By Theorem 3, this is equivalent to the Fourier extension conjecture, and this completes the proof of Theorem 10. ∎
@@ 4. Appendix: Application to the probabilistic Fourier extension theorem
taken over all $s_{1}\leq s_{2}$ and $s_{2},s_{3}\in\left(\left(1-\delta\right)s,\left(1+\delta\right)s\right)$, all $f_{k}\in L^{q}\left(U_{k}\right)$, and all triples $\left(U_{1},U_{2},U_{3}\right)\subset U^{3}$ that satisfy the weak $\nu$-disjoint condition,
@@
It should be possible to prove this square function variant by tracing through the proof of Theorem 10 above and making modifications for the square function. See for example [RiSa2], where these modifications are carried out for ‘father’ wavelets $\varphi_{I}$ smoothly adapted to a square II, and their projections $\bigtriangleup_{I}f\equiv\left\langle f,\varphi_{I}\right\rangle\varphi_{I}$. However, an additional difficulty arises for smooth Alpert wavelets $h_{I;\kappa}^{\eta}$, since the pseudoprojections $\bigtriangleup_{I;\kappa}^{\eta}f\equiv\left\langle S_{\kappa,\eta}^{-1}f,h_{I;\kappa}\varphi_{I}\right\rangle h_{I;\kappa}^{\eta}$ involve an operator $S_{\kappa,\eta}^{-1}$ that is bounded only on $L^{p}$ for $1<p<\infty$, and not on $L^{\infty}$. In this case, parabolic rescalings must be performed with $f\in L^{p}\left(U\right)$, which introduces an additional growth factor $C_{p}2^{\frac{2}{p}s}$, that turns out to be harmless because one can take $p$ arbitrarily large. As mentioned earlier, the proof of this conjecture will be addressed in a future paper.